OpenStudy (perfecta):

The value of acceleration due to gravity (g) on Pluto is about 0.61 meters/second2. How much will an object that weighs 250 newtons on Earth weigh on Pluto? Note that the value of acceleration due to gravity on Earth is 9.8 meters/second2.

OpenStudy (chrishogan):

F=ma 1 (kg)=9.8 (N) so 250 (N)=25.5 (kg) then, in Pluto F=ma=25.5×0.61=15.56(N)

OpenStudy (chrishogan):

Did this help you? @PerfectA

OpenStudy (perfecta):

Yes, Thanks @ChrisHogan

OpenStudy (chrishogan):

Glad to help. :)